OpenStudy (destinyyyy):

A chemist needs 190 milliliters of a 47% solution but has only 25% and 63% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.

OpenStudy (perl):

If we assume that x = number of milliliters of 25% solution y = number of milliliters of 63% solution x + y = 190 .25x + .63y = 190 * .47

OpenStudy (destinyyyy):

Okay

OpenStudy (destinyyyy):

I match up the y's to cancel them out correct?

OpenStudy (perl):

yes, you can solve by substitution or elimination

OpenStudy (destinyyyy):

I got x=317.38 which cant be correct

OpenStudy (perl):

x + y = 190 .25x + .63y = 190 * .47 multiply the second equation by 100. that will get rid of decimals. x+y=190 25x + 63y = 8930

OpenStudy (destinyyyy):

x=-230 then I plug this into x+y=190

OpenStudy (perl):

x should be positive

OpenStudy (destinyyyy):

Yes ... For x+y=190 times it all by -63 to cancel out the ys ... But then I get -38x=8740

OpenStudy (perl):

Did you get this? -63x -63y = -11970 25x + 63y = 8930

OpenStudy (perl):

I multiplied first equation by -63

OpenStudy (destinyyyy):

-.- oh my gosh i left it as -190

OpenStudy (perl):

:D

OpenStudy (destinyyyy):

x=80 and y=110

OpenStudy (destinyyyy):

Thank you!